OpenStudy (anonymous):

Write the inverse variation equation for the relationship: y varies inversely with x and y = 4 when x = 2.

OpenStudy (lgbasallote):

\[\large \begin{align} y & = \frac{k}{x}\\\\ 4 & = \frac{k}{2}\\\\ 8 & = k \end{align}\]

OpenStudy (lgbasallote):

shouldnt there be a second situatioN?

OpenStudy (anonymous):

well yes, but I already got the main answer to plug in. thanks!

OpenStudy (anonymous):

and also, technically thats yx=k right? easier way to write it?

OpenStudy (lgbasallote):

yes